OpenStudy (anonymous):

another word problem two bars togeather make 550,000,000 but bar one makes 450,000 more than the other. how much does each bar make

OpenStudy (anonymous):

sorry for the misspell

OpenStudy (anonymous):

\[x+y=550,000,000\] \[x-y=450,000\] your math teacher must hate you

OpenStudy (lgbasallote):

let x = bar 1 y = bar 2 it says bar one makes 450,000 more than the other so x = 450,000 + y two bars together make 550,000,000 so x + y = 550,000,000 since x = 450,000 + y just substitute 450,000 + y + y = 550,000,000 solve for y does that help?

OpenStudy (anonymous):

for some reason I just do not get word problems. I mean how to set them up so that i can work them

OpenStudy (anonymous):

since you are going to have to use a calculator anyways, just do this http://www.wolframalpha.com/input/?i=x%2By%3D550%2C000%2C000%2Cx-y%3D450%2C000